OpenStudy (anonymous):

find an equation of variation where y varies directly as the square of x, and y=54 when x=3 √2

OpenStudy (shamil98):

direct variation is x1/x2 = y1/y2

OpenStudy (anonymous):

" y varies directly as the square of x" means y = cx^2 for some constant c (you need to figure out c)

OpenStudy (anonymous):

Do you know how to figure out what c is?